Oxygen gas reacts with powdered aluminum according to the reaction:
4Al(s)+3O2(g)→2Al2O3(s)

Part A

What volume of O2 gas (in L), measured at 762  Torr and 32  ∘C , completely reacts with 52.1  g of Al?
Answer #1

Molar mass of Al = 26.98 g/mol

mass of Al = 52.1 g

mol of Al = (mass)/(molar mass)

= 52.1/26.98

= 1.931 mol

According to balanced equation

mol of O2 reacted = (3/4)* moles of Al

= (3/4)*1.931

= 1.448 mol

Given:

P = 762.0 torr

= (762.0/760) atm

= 1.0026 atm

n = 1.448 mol

T = 32.0 oC

= (32.0+273) K

= 305 K

use:

P * V = n*R*T

1.0026 atm * V = 1.448 mol* 0.08206 atm.L/mol.K * 305 K

V = 36.1459 L

Answer: 36.1 L